# Conan-embedding: General Text Embedding with More and Better Negative Samples

### v1

@@ 2.1 Training Workflow
$\mathcal{L}_{\text{neg}}=-\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(x_{i},x_{i}^{+}))}{\sum_{j=1}^{N}\exp(\text{sim}(x_{i},x_{j}))}$ (1)
@@
$\mathcal{L}_{\text{cos}}=\log\left(1+\sum\limits_{\text{sim}(i,j)>\text{sim}(k,l)}e^{\frac{\lambda(\cos(u_{k},u_{l})-\cos(u_{i},u_{j}))}{\tau}}\right)$ (2)
where $\tau$ is the scale temperature, $\cos(\cdot)$ is the cosine similarity function, and $s(u,v)$ is the similarity between $u$ and $v$.
@@ 2.3 Cross-GPU Batch Balance Loss
To better leverage hard examples, we adopted the Cross-GPU Batch Balance Loss (CBB). Previous approaches Li et al. (2023b) typically assign a task to each batch randomly during the training process. For example, in iteration 0, samples from the STS task are selected, and the corresponding STS loss is used for backward to obtain gradients and update weights. In iteration 1, the Retri task or the CLS task might be assigned. We refer to this as sequential random task training. Such training almost invariably leads to inconsistencies between the optimization search space of a single iteration and the global optimization search space of the embedding model, resulting in oscillations during the training process and the inability to achieve a global optimum. We demonstrate this phenomenon in subsequent analysis.
@@
$\mathcal{L}_{\text{CBB}}=-\frac{1}{n}\sum_{i}\log\frac{e^{s(x,y^{+})/\tau}}{e^{s(x_{i},y_{i}^{+})/\tau}+\sum_{k=1}^{N}\sum_{j=1}^{n}e^{s(x_{i},y_{j}^{-})/\tau}}+\beta\times\mathcal{L}_{\text{cos}}$ (3)
where $s(x,y^{+})$ is a scoring function between query $q$ and positive text $y$, often defined as the cosine similarity, $N$ is the number of GPUs sharing the query $q$ and positive text $y$, and $\tau$ is the scale temperature. We set $\beta$ to 0.8 empirically.

### v2

@@ 2.1.1 Pre-training
$\mathcal{L}_{neg}=-\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(x_{i},y_{i}^{+}))}{\sum_{j=1}^{M}\exp(\text{sim}(x_{i},y_{i}))}$ (1)
$x_{i}$ represents the query of the positive sample, $y_{i}^{+}$ represents the passage of the positive sample, and $y_{i}$ represents the passages of other samples in the same batch, which are considered as negative samples.
@@ 2.1.2 Supervised Fine-tuning
$\mathcal{L}_{cos}=\log\left(1+\sum\limits_{\text{sim}(i,j)>\text{sim}(k,l)}\exp\left(\frac{\cos(x_{k},x_{l})-\cos(x_{i},x_{j})}{\tau}\right)\right)$ (2)
where $\tau$ is the scale temperature, $\cos(\cdot)$ is the cosine similarity function, and $sim(k,l)$ is the similarity between $x_{i}$ and $x_{j}$.
@@ 2.3 Cross-GPU Batch Balance Loss
To better leverage hard examples, we adopted the Cross-GPU Batch Balance Loss (CBB). Previous approaches Li et al. (2023b) typically assign a task to each batch randomly during the training process. For example, in iteration 0, samples from the STS task are selected, and the corresponding STS loss is used for backward to obtain gradients and update weights. In iteration 1, the Retri task might be assigned. We refer to this as sequential random task training. Such training often results in inconsistencies between the search space optimized in a single iteration and the global search space of the embedding model, which can cause oscillations during training. These oscillations hinder the model’s ability to converge to a global optimum, making it more challenging to achieve the best possible performance. We demonstrate this phenomenon in Sec. 3.5.
@@
$\mathcal{L}_{\text{CBB}}=-\frac{1}{n}\sum_{i}\log\frac{\exp(s(x_{i},y_{i}^{+})/\tau)}{\exp(s(x_{i},y_{i}^{+})/\tau)+\sum_{k=1}^{N}\sum_{j=1}^{n}\exp(s(x_{i},y_{j}^{-})/\tau)}+\beta\times\mathcal{L}_{\text{cos}}$ (3)
where $s(x_{i},y_{i}^{+})$ is a scoring function between query $x_{i}$ and positive text $y_{i}^{+}$, often defined as the cosine similarity, $N$ is the number of GPUs sharing the query $x_{i}$ and positive text $y_{i}^{+}$, and $\tau$ is the scale temperature. We set $\beta$ to 0.8 empirically.